OpenStudy (anonymous):

Express the inverse of the following matrix (assuming it exists) as a matrix containing expressions in terms of k. -3 0 k 12 8 16 4 2 4

OpenStudy (zzr0ck3r):

um you gave a list of numbers that in no way resembles a matrix

OpenStudy (anonymous):

\[\left[\begin{matrix}3 & 0 & k \\ 12 & 8 & 16\\ 4 & 2 & 4\end{matrix}\right]^{-1}\]

OpenStudy (zzr0ck3r):

can you reduce it to I?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

i write where i am write now

OpenStudy (zzr0ck3r):

k brb wife just got home.

OpenStudy (anonymous):

this is just the coefficient matrix\[\left[\begin{matrix}1 & 0 & 0 \\ 0 & 1 & 2\\ 0 & 0 & k\end{matrix}\right]\]

OpenStudy (anonymous):

Use Gauss-Jordan elimination on \[\left[\begin{matrix}3 & 0 & k & 1 & 0 & 0\\ 12 & 8 & 16 & 0 & 1 & 0\\ 4 & 2 & 4 & 0 & 0 & 1\end{matrix}\right]\] That's how I do it. (There's probably an easier way..)

OpenStudy (anonymous):

yes i started w/ that

OpenStudy (anonymous):

Hmm, is that 3 in a_11 positive or negative?

OpenStudy (anonymous):

this is where im at now\[\left[\begin{matrix}1 & 0 & 0 & 0 & \frac{ -1 }{ 4} & 1 \\ 0 & 1 & 2 & 0 & \frac{ 1 }{ 2 } & \frac{ -3 }{ 2 } \\ 0 & 0 & k & 1 & \frac{ -3 }{ 4 } & 3\end{matrix}\right]\]

OpenStudy (anonymous):

postive

OpenStudy (anonymous):

how would i express k on the inverse matrix

OpenStudy (anonymous):

Continue to reduce it using row operations until you have only the identity matrix on the left.

OpenStudy (anonymous):

could i multiplay R3 by 1/k?

OpenStudy (anonymous):

multiply*

OpenStudy (anonymous):

Looks like you have to.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The last step ought to be subtracting your new R3 from R2.

OpenStudy (anonymous):

*Sorry, twice R3 from R2..